OpenStudy (anonymous):

Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution.

OpenStudy (magbak):

@GoatsCheese

OpenStudy (anonymous):

\[\frac{ x+a }{ ax }=\frac{ b }{ x }\]

OpenStudy (anonymous):

@ninjasandtigers

OpenStudy (anonymous):

pick a number for a and b

OpenStudy (anonymous):

2 and 4